OpenStudy (anonymous):

is 5000 a power of 10

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

No. 1, 10, 100, 1000, 10000, 100000, ... are all powers of 10, as are \(\frac{1}{10}, \frac{1}{100}, ...\). But 5000 is only a multiple of 10. It is equivalent to \(10^3 \times 5\).